# Efficient Construction of Model Family through Progressive Training Using Model Expansion

### v1

@@ Abstract
As Large Language Models (LLMs) gain widespread practical application, providing the model family of different parameter sizes has become standard practice to address diverse computational requirements. Conventionally, each model in a family is trained independently, resulting in computational costs that scale additively with the number of models. We propose an efficient method for constructing the model family through progressive training, where smaller models are incrementally expanded to larger sizes to create a complete model family. Through extensive experiments with a model family ranging from 1B to 8B parameters, we demonstrate that our method reduces computational costs by approximately 25% while maintaining comparable performance to independently trained models. Furthermore, by strategically adjusting maximum learning rates based on model size, our method outperforms the independent training across various metrics. Beyond performance gains, our approach offers an additional advantage: models in our family tend to yield more consistent behavior across different model sizes.
@@ 1 Introduction
As Large Language Models (LLMs) gain widespread practical application, providing models with (i) a consistent architecture and (ii) different parameter sizes (hereafter referred to as a model family) has become standard practice in the NLP community. For example, Llama 3.1 consists of models with 8B, 70B, and 405B parameters (AI@Meta, 2024), and Gemma 3 includes 1B, 4B, 12B and 27B parameters (GemmaTeam, 2025) . Similarly, Qwen2.5 provides a model family with 0.5B, 1.5B, 3B, 7B, 14B, 32B and 72B parameters (Yang et al., 2024b).
@@
Offering a model family is a typical approach to addressing diverse computational requirements and use cases. Smaller models enable efficient processing and rapid response for daily tasks while facilitating deployment in environments with strict computational requirements, such as smartphones and edge devices (Abdin et al., 2024). Conversely, larger models are deployed for scenarios requiring advanced reasoning capabilities and complex task processing, typically on large-scale servers (Wei et al., 2022).
@@
The conventional approach to constructing a model family involves training each model independently from scratch. However, training large-scale models requires substantial resources, e.g., thousands of GPU days (Touvron et al., 2023). The total computational cost of constructing a model family poses a significant burden on its builders. (Figure 1 (Top)). This fact motivates us to explore more efficient methods for constructing a model family.
@@
We identify model expansion as a potential approach to more efficiently constructing model families, including large-scale models. Model expansion is a method of leveraging the parameters of pre-trained smaller models as an effective initialization for larger models (Chen et al., 2022; Du et al., 2024). However, existing work on model expansion typically focuses on producing a single final model, paying less attention to the potential utility of intermediate models. In this work, we empirically explore model expansion as a means to efficiently construct a model family. Specifically, we propose a method that repeatedly applies model expansion to construct models from smaller to larger sizes, namely progressive training, thereby reducing the total training cost of constructing a model family (Figure 1 (Bottom)).
@@
Through our experiments with a model family ranging from 1B to 8B sizes, we demonstrate that our proposed method reduces the total computational cost of constructing a complete model family by approximately 25% compared to training each model independently111 This 25% reduction translates to savings of about 3.2K GPU hours in our experimental setup. . Furthermore, by adjusting the maximum learning rate according to model size, our method achieves performance that consistently outperforms independently trained models across various benchmark tasks. Additionally, we show that our progressive training approach produces models with greater behavioral consistency across the model family, as evidenced by lower Kullback-Leibler (KL) divergence between models’ probability distributions.
@@
Figure 1: Comparison of approaches for constructing a model family. (Top): The conventional approach where each model in the family (2B, 4B, 8B) is trained independently from scratch. The total computational cost equals the sum of individual costs for all models. (Bottom): The proposed progressive training utilizes model expansion, where smaller models are expanded to initialize larger ones. The total cost equals only that of the largest model (8B).
@@ 2 Task Definition and Notation Rules
We consider a model family consisting of models with monotonically increasing sizes and consistent architecture. Formally, we define this model family as a sequence of models of sizes $[X_{1},X_{2},\ldots,X_{n}]$, corresponding sequence of model parameters $[\theta_{1},\theta_{2},\ldots,\theta_{n}]$, where $\theta_{i}\in\mathbb{R}^{X_{i}}$ represents the parameters of the $i$-th model. For each model size $X_{i}$, let $T_{i}$ denote the number of tokens used to train the corresponding model. The computational cost (FLOPs) required to train a given model is determined by the model size $X_{i}$ and the number of tokens $T_{i}$. Following Brown et al. (2020), we define FLOPs as: $\mathrm{FLOPs}(X_{i},T_{i})=6X_{i}T_{i}$. Note that our analysis does not rely on this specific calculation and can be generalized to other reasonable cost approximations (Hoffmann et al., 2022).
Let $T_{i}^{\textrm{scratch}}$ denote the number of tokens for training a randomly initialized model with size $X_{i}$ from scratch. Then, the total computational cost of constructing an entire model family is $\sum_{i=1}^{n}\mathrm{FLOPs}(X_{i},T_{i}^{\textrm{scratch}})$. The task of this research is to reduce this computational cost and maintain the performance of each model. Formally, we construct each model by satisfying the following constraint:
@@
where $T_{i}^{\mathrm{prog}}$ is the number of tokens used for the training model of size $X_{i}$ in our progressive training approach (Section 3).
@@ 3 Method: Progressive Training
We propose a method for constructing language models of multiple sizes more efficiently than the common practice, i.e., training each model independently from scratch. In the common approach, the cost of training a model of size $X_{i}$ from scratch using $T_{i}^{\textrm{scratch}}$ tokens222 Note that the choice of $T_{i}^{\textrm{scratch}}$ is arbitrary depending on available resources, but most researchers adopt the Chinchilla law (Hoffmann et al., 2022) as a principled approach to determine the compute-optimal number of training tokens. Our experimental setup is also based on the Chinchilla law (Section 4.2). is $\mathrm{FLOPs}(X_{i},T_{i}^{\textrm{scratch}})$, resulting in the total cost of constructing a model family $\sum_{i=1}^{n}\mathrm{FLOPs}(X_{i},T_{i}^{\textrm{scratch}})$. A straightforward approach for satisfying Equation (1) is to reduce the number of training tokens for each model by choosing $T_{i}^{\textrm{prog}}$ such that $T_{i}^{\textrm{scratch}}>T_{i}^{\textrm{prog}}$ for $i>1$. However, a naive reduction of training tokens will lead to performance degradation if other conditions do not change. Instead, when constructing a model family, we jumpstart the training of $i$-th model of size $X_{i}$ for $i>1$ by using a better initialization than random, which is obtained by applying model expansion to its preceding smaller model of size $X_{i-1}$. We call this procedure progressive training.
Specifically, our progressive training starts with training an initial model of size $X_{1}$ with $T_{1}^{\textrm{scratch}}$ tokens. In each subsequent stage, we initialize the next model using a model expansion method $f$,
@@
where $f(\cdot;X_{i+1}):\mathbb{R}^{X_{i}}\to\mathbb{R}^{X_{i+1}}$ is an off-the-shelf model expansion method. The expanded parameter $\theta_{i+1}^{\textrm{init}}$ serves as an effective initialization for the subsequent model training.
There is a vast range of possible token allocation patterns $[T_{1}^{\textrm{prog}},\ T_{2}^{\textrm{prog}},...,T_{n}^{\textrm{prog}}]$ that satisfy Equation (1)’s condition regarding the total computational cost. In this work, we determine each $T_{i+1}^{\textrm{prog}}$, ensuring that the total computational cost equals that of training the largest model $X_{n}$ from scratch:
@@
Specifically, we set $T_{i+1}^{\textrm{prog}}$ so that it satisfies the following FLOPs constraint:
@@
Intuitively, we allocate the computational cost of training a model of size $X_{i+1}$ from scratch, subtracting the computational cost already used in previous stages. Through this procedure, we can obtain an entire model family from size $X_{1}$ to $X_{n}$ with the total computational cost, which equals to $\mathrm{FLOPs}(X_{n},T_{n}^{\textrm{scratch}})$.
@@
In this work, we employ bert2BERT (Chen et al., 2022) as model expansion method for the following reason. Unlike approaches that focus solely on depth expansion (such as stacking (Du et al., 2024)), bert2BERT allows us to increase both the width and depth dimensions of Transformer models (Vaswani et al., 2017) simultaneously, providing greater flexibility333 While recent methods like LEMON (Wang et al., 2024b) have emerged, LEMON builds upon bert2BERT’s foundations with relatively minor modifications to the core mechanisms. We adopt bert2BERT for its simplicity and ease of implementation. . Specifically, for width expansion, bert2BERT duplicates linear layer weights to increase hidden dimensions, while for depth expansion, it duplicates pre-trained layers and stacks them [^4] The original paper of bert2BERT (Chen et al., 2022) proposes two variants for model expansion: AKI and FPI. We exclusively use AKI in this work.. This flexibility to expand in multiple dimensions is particularly valuable as we repeatedly apply model expansion in progressive training.
@@ 4 Experiment
To demonstrate the effectiveness of progressive training, we pre-train and compare two model families: one with independently training each model from scratch (Independent) and the other with progressive training (Progressive). Specifically, we show that progressive training improves computational efficiency over independent training and the performance of the two model families is comparable.
@@ 4.1 Experimental Setup
We used FineWeb-Edu (Penedo et al., 2024) as the training data for pre-training. We adopted GPT-2 (Radford et al., 2019) tokenizer for tokenization. While progressive training can accommodate any parameter size increase, in this work, we adopted a setting where the parameter count doubles at each stage. Specifically, we construct a model family consisting of $[X_{1}=1\textrm{B},X_{2}=2\textrm{B},X_{3}=4\textrm{B},X_{4}=8\textrm{B}]$555 The specific width and depth of each model are described in Appendix A. . Our models follow the Llama architecture (Touvron et al., 2023). The maximum sequence length for inputs was set to 1024. Furthermore, each model was trained with a cosine learning rate scheduler, using a maximum learning rate of $3.0\times 10^{-4}$.
For the evaluation of each pre-trained model in the family, we used perplexity on the validation data (Valid) from FineWeb-Edu and WikiText (Merity et al., 2017). Additionally, to comprehensively assess the performance of pre-trained models, we measured zero-shot performance across multiple downstream tasks. Specifically, we adopted tasks including language modeling (LAMBADA (Paperno et al., 2016)), commonsense reasoning (WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019)), and question answering (ARC-e, ARC-c (Clark et al., 2018), OBQA (Mihaylov et al., 2018)).
@@ 4.2 Training Data Size and Computational Cost
In the progressive training approach, we determine the number of training tokens $T_{i}^{\textrm{prog}}$ for each model to satisfy the constraint defined in Equation (4), ensuring that the total computational cost equals that of training the largest model from scratch. For our 2x Chinchilla law setting, this computational requirement resulted in progressive training with the following tokens: $T_{1}^{\textrm{prog}}=40\textrm{B},T_{2}^{\textrm{prog}}=60\textrm{B},T_{3}^{\textrm{prog}}=120\textrm{B},T_{4}^{\textrm{prog}}=240\textrm{B}$666 For the Chinchilla law setting, $T_{i}^{\textrm{prog}}$ is shown in Appendix B..
One of the key advantages of progressive training is the ability to obtain models of multiple sizes efficiently. As defined in Section 2, the computational cost is calculated as $\textrm{FLOPs}=6XT$, where $X$ is the parameter count of each model and $T$ is the number of training tokens. For example, independently training each model in the family $[1\mathrm{B},2\mathrm{B},4\mathrm{B},8\mathrm{B}]$ under the 2x Chinchilla law would require FLOPs of $[0.24\textrm{Z},0.96\textrm{Z},3.84\textrm{Z},15.4\textrm{Z}]$ respectively. Thus, constructing the entire model family would require $20.4\mathrm{ZFLOPs}$ in total.
In contrast, with progressive training, the computational cost required to construct the entire model family is equivalent to the computational cost required to train the largest-size, i.e., 8B, model. Therefore, the computational cost is $15.4\mathrm{ZFLOPs}$. Compared to training each model independently, progressive training can reduce computational cost by approximately 25%. This demonstrates that progressive training improves the computational efficiency of constructing a model family.
@@ 4.3 Results
Table 3: Evaluation results of pre-trained models with maximum learning rate adjustment strategy described in Section 4.4. Progressive models consistently outperform Independent models. +Fixed Data indicates training on a fixed dataset of 320B tokens.
@@
Tables 1 and 2 present the evaluation results of pre-trained models in the Chinchilla law and 2x Chinchilla law settings, respectively. Our progressive training approach (Progressive) achieves performance comparable to or better than models that are independently trained from scratch (Independent) across different parameter sizes in both settings. It is particularly noteworthy that, for the largest model (8B), Progressive shows improvements in perplexity and most downstream tasks.
@@
In Figure 2 (Left), for a given computational cost, i.e., ZFLOPs, Progressive consistently outperforms Independent. In addition, Progressive reaches the training loss of Independent while achieving a 26% reduction in FLOPs.
@@
[…]
The result is shown in Table 2, in which 8B Progressive+Fixed Data achieves comparable performance. Thus, we confirm that the performance improvements from progressive training do not depend on the addition of more unique data.

### v2

@@ Abstract
As Large Language Models (LLMs) gain widespread practical application, offering model families with varying parameter sizes has become standard practice to accommodate diverse computational requirements. Traditionally, each model in the family is trained independently, incurring computational costs that scale additively with the number of models. In this work, we propose an efficient method for constructing model families via progressive training, where smaller models are incrementally expanded to larger sizes to create a complete model family. Through extensive experiments on a model family ranging from 1B to 8B parameters, we show that our approach reduces total computational cost by approximately 25% while maintaining comparable performance to independently trained models. Moreover, by strategically adjusting the maximum learning rate based on model size, our method outperforms the independent training across various metrics. Beyond these improvements, our approach also fosters greater consistency in behavior across model sizes.
@@ 1 Introduction
As Large Language Models (LLMs) gain widespread practical application, providing models with (i) a consistent architecture and (ii) varying parameter sizes (hereafter referred to as a model family) has become standard practice in the NLP community. For instance, Llama 3.1 includes models with 8B, 70B, and 405B parameters (AI@Meta, 2024), while Gemma 3 provides 1B, 4B, 12B, and 27B parameter variants (GemmaTeam, 2025) . Similarly, Qwen2.5 offers a model family with 0.5B, 1.5B, 3B, 7B, 14B, 32B, and 72B parameters (Yang et al., 2024b).
@@
Such model families are designed to address a wide range of computational constraints and application scenarios. Smaller models offer faster inference and lower resource consumption, making them suitable for daily tasks and deployment in resource-constrained environments such as smartphones and edge devices (Abdin et al., 2024). In contrast, larger models are deployed for scenarios requiring advanced reasoning capabilities and complex task processing, typically on large-scale servers (Wei et al., 2022).
@@
The standard approach to constructing a model family involves training each model independently from scratch. However, training large-scale models demands extensive resources, e.g., thousands of GPU days (Touvron et al., 2023). The total computational cost of constructing a model family poses a significant burden on its builders. (Figure 1 (Top)). This motivates us to explore more efficient methods for model family construction.
@@
We identify model expansion as a potential approach to more efficiently constructing model families, including large-scale models. Model expansion leverages the parameters of pre-trained smaller models as initialization for training larger models (Chen et al., 2022; Du et al., 2024). However, prior work on model expansion has primarily aimed at producing a single final model, with limited focus on the potential utility of intermediate models.
In this work, we empirically explore model expansion as a means to efficiently construct a model family. Specifically, we propose a method that repeatedly applies model expansion to construct models from smaller to larger sizes, namely progressive training, thereby reducing the total training cost of constructing a model family (Figure 1 (Bottom)).
@@
Through our experiments with a model family ranging from 1B to 8B parameters, we demonstrate that the proposed method reduces the total computational cost of constructing a complete model family by approximately 25% compared to training each model independently [^1] This 25% reduction corresponds to savings of approximately 3.2K GPU hours in our experimental setup. . Moreover, by adjusting the maximum learning rate based on model size, our method consistently achieves superior performance across a range of benchmark tasks relative to independently trained models. We also found that progressive training yields greater behavioral consistency across the model family, as indicated by lower Kullback-Leibler (KL) divergence between models’ output distributions.
@@
Figure 1: Comparison of approaches for constructing a model family. (Top): Conventional approach, where each model in the family (2B, 4B, 8B) is trained independently from scratch. The total computational cost is the sum of the training costs for all individual models. (Bottom): Proposed progressive training utilizes model expansion, where smaller models are expanded to initialize larger ones. The total cost equals only that of the largest model (8B).
@@ 2 Task Definition and Notation Rules
We consider a model family consisting of models with monotonically increasing sizes and consistent architecture. Formally, we define this model family as a sequence of models of sizes $X_{1},X_{2},\ldots,X_{n}]$ and corresponding sequence of model parameters $\theta_{1},\theta_{2},\ldots,\theta_{n}]$, where $\theta_{i}\in\mathbb{R}^{X_{i}}$ denotes the parameters of the $i$-th model. For each model of size $X_{i}$, let $T_{i}$ represent the number of training tokens used to train the corresponding model. The computational cost (FLOPs) required to train a given model is determined by the model size $X_{i}$ and the number of tokens $T_{i}$. Following Brown et al. (2020), we define FLOPs as: $\mathrm{FLOPs}(X_{i},T_{i})=6X_{i}T_{i}$. Note that our analysis does not depend on this specific formulation and can be generalized to other reasonable cost approximations (Hoffmann et al., 2022).
Let $T_{i}^{\textrm{scratch}}$ denote the number of tokens for training a randomly initialized model of size $X_{i}$ from scratch. The total computational cost of constructing the entire model family is then given by $\sum_{i=1}^{n}\mathrm{FLOPs}(X_{i},T_{i}^{\textrm{scratch}})$. The task of this work is to reduce the overall computational cost while preserving the performance of each model. Formally, we aim to construct each model under the following constraint:
@@
where $T_{i}^{\mathrm{prog}}$ denotes the number of tokens used to train the model of size $X_{i}$ under the progressive training approach (Section 3).
@@ 3 Method: Progressive Training
We propose a method for constructing language models of varying sizes more efficiently than the common practice, i.e., training each model independently from scratch. In the common approach, the cost of training a model of size $X_{i}$ from scratch using $T_{i}^{\textrm{scratch}}$ tokens [^2] Note that the choice of $T_{i}^{\textrm{scratch}}$ is arbitrary depending on available resources, though most researchers follow the Chinchilla law (Hoffmann et al., 2022) as a principled approach to determine the compute-optimal number of training tokens. Our experimental setup also adopts the Chinchilla law (Section 4.2). is $\mathrm{FLOPs}(X_{i},T_{i}^{\textrm{scratch}})$, leading to a total cost of $\sum_{i=1}^{n}\mathrm{FLOPs}(X_{i},T_{i}^{\textrm{scratch}})$ for the entire model family. A simple strategy to satisfy Equation (1) is to reduce the number of training tokens for each model by selecting $T_{i}^{\textrm{prog}}$ such that $T_{i}^{\textrm{scratch}}>T_{i}^{\textrm{prog}}$ for $i>1$. However, naively reducing the number of training tokens typically results in degraded performance if other factors remain unchanged. Instead, we initialize the $i$-th model of size $X_{i}$ for $i>1$ by using a better initialization than random, which is obtained by applying model expansion to the previously trained model sizesize $X_{i-1}$. We refer to this process as progressive training.
Specifically, our progressive training begins by training an initial model of size $X_{1}$ using $T_{1}^{\textrm{scratch}}$ tokens. At each subsequent stage, we initialize the next model using a model expansion method $f$,
@@
where $f(\cdot;X_{i+1}):\mathbb{R}^{X_{i}}\to\mathbb{R}^{X_{i+1}}$ is an off-the-shelf model expansion method. The expanded parameters $\theta_{i+1}^{\textrm{init}}$ serve as an effective initialization for training the subsequent model.
There exists a wide range of possible token allocation patterns $T_{1}^{\textrm{prog}},\ T_{2}^{\textrm{prog}},...,T_{n}^{\textrm{prog}}]$ that satisfy the constraint in Equation (1)’s condition regarding the total computational cost. In this work, we determine each $T_{i+1}^{\textrm{prog}}$ such that the overall computational cost matches that of training the largest model $X_{n}$ from scratch:
@@
Specifically, we set $T_{i+1}^{\textrm{prog}}$ to satisfy the following FLOPs constraint:
@@
Intuitively, we allocate the computational cost of training a model of size $X_{i+1}$ from scratch, subtracting the computational cost already used in previous stages. Through this procedure, we can obtain an entire model family from size $X_{1}$ to $X_{n}$ with a total computational cost equal $\mathrm{FLOPs}(X_{n},T_{n}^{\textrm{scratch}})$.
@@
In this work, we adopt bert2BERT (Chen et al., 2022) as our model expansion method for the following reason. Unlike approaches that focus solely on depth expansion (such as stacking (Du et al., 2024)), bert2BERT allows us to increase both the width and depth dimensions of Transformer models (Vaswani et al., 2017), offering greater flexibility [^3] Although recent methods such as LEMON (Wang et al., 2024b) have emerged, LEMON builds upon the foundation of bert2BERT with only minor modifications to its core mechanisms. We chose bert2BERT for its simplicity and ease of implementation. . Specifically, for width expansion, bert2BERT increases the hidden dimensions by duplicating the weights of linear layers; for depth expansion, it stacks additional layers by duplicating pre-trained ones [^4] The original bert2BERT paper (Chen et al., 2022) introduces two expansion variants: AKI and FPI. In this work, we exclusively adopt the AKI. This flexibility to expand along both dimensions is particularly valuable as we repeatedly apply model expansion in progressive training.
@@ 4 Experiment
To evaluate the effectiveness of progressive training, we pre-train and compare two model families: one with independently training each model from scratch (Independent) and the other with progressive training (Progressive). Specifically, we demonstrate that progressive training improves computational efficiency over independent training, and the performance of the two model families is comparable.
@@ 4.1 Experimental Setup
We used FineWeb-Edu (Penedo et al., 2024) as the training data for pre-training and adopted the GPT-2 (Radford et al., 2019) tokenizer for tokenization. While progressive training can accommodate any parameter size increase, in this work, we adopted a configuration where the parameter count doubles at each stage. Specifically, we construct a model family with sizes $X_{1}=1\textrm{B},X_{2}=2\textrm{B},X_{3}=4\textrm{B},X_{4}=8\textrm{B}]$ [^5] The specific width and depth settings for each model are detailed in Appendix A. . All models follow the Llama architecture (Touvron et al., 2023) with a maximum input sequence of 1024 tokens. Each model was trained using a cosine learning rate scheduler, with a maximum learning rate of $3.0\times 10^{-4}$.
To evaluate each pre-trained model in the family, we measured perplexity on the validation data (Valid) from FineWeb-Edu and WikiText (Merity et al., 2017). To more comprehensively assess the performance of pre-trained models, we also evaluated zero-shot performance across multiple downstream tasks. Specifically, we include tasks including language modeling (LAMBADA (Paperno et al., 2016)), commonsense reasoning (WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019)), and question answering (ARC-e, ARC-c (Clark et al., 2018), OBQA (Mihaylov et al., 2018)).
@@ 4.2 Training Data Size and Computational Cost
In the progressive training approach, we determine the number of training tokens $T_{i}^{\textrm{prog}}$ for each model to satisfy the constraint defined in Equation (4), ensuring it matches the total computational cost equals that of training the largest model from scratch. Under the 2x Chinchilla law setting, this constraint yields the following token allocations: $T_{1}^{\textrm{prog}}=40\textrm{B},T_{2}^{\textrm{prog}}=60\textrm{B},T_{3}^{\textrm{prog}}=120\textrm{B},T_{4}^{\textrm{prog}}=240\textrm{B}$ [^6] For the Chinchilla law setting, $T_{i}^{\textrm{prog}}$ is shown in Appendix B..
One of the key advantages of progressive training is its ability to obtain models of multiple sizes efficiently. As defined in Section 2, the computational cost is calculated as $\textrm{FLOPs}=6XT$, where $X$ denotes the number of model parameters and $T$ is the number of training tokens. For instance, independently training each model in the family $1\mathrm{B},2\mathrm{B},4\mathrm{B},8\mathrm{B}]$ under the 2x Chinchilla law would require FLOPs of $0.24\textrm{Z},0.96\textrm{Z},3.84\textrm{Z},15.4\textrm{Z}]$ respectively. Thus, constructing the entire model family would require $20.4\mathrm{ZFLOPs}$ in total.
In contrast, with progressive training, the total computational cost required to construct the entire model family is equivalent to the computational cost required to train the largest-size, i.e., 8B, model. Therefore, the computational cost is $15.4\mathrm{ZFLOPs}$. Compared to training each model independently, progressive training can reduce computational cost by approximately 25%. This demonstrates that progressive training improves the computational efficiency of constructing a model family.
@@ 4.3 Results
Table 3: Evaluation results of pre-trained models with the maximum learning rate adjustment strategy described in Section $4.$ Progressive models consistently outperform Independent models. +Fixed Data indicates training on a fixed dataset of 320B tokens.
@@
Tables 1 and 2 present the evaluation results of pre-trained models in the Chinchilla law and 2x Chinchilla law settings, respectively. Our progressive training approach (Progressive) achieves performance comparable to or better than models that are independently trained from scratch (Independent) across different parameter sizes in both settings. Notably, for the largest model (8B), Progressive shows improvements in perplexity and most downstream tasks.
@@
In Figure 2 (Left), for a given computational budget, i.e., ZFLOPs, Progressive consistently outperforms Independent. In addition, Progressive reaches the same training loss as Independent while reducing the total FLOPS by 26%.
@@
For each model size $X_{i}(i>1)$ in our model family, we use fewer tokens in progressive training than independent training ($T_{i}^{\textrm{prog}}<T_{i}^{\textrm{scratch}}$). However, when considering the entire model family construction process, progressive training consumes more total tokens than training only the largest model from scratch. As detailed in Section 4.2, the total number of tokens processed in progressive training is $\sum_{i=1}^{4}{T_{i}^{\textrm{prog}}}=460\textrm{B}$ tokens, which exceeds the $T_{4}^{\textrm{scratch}}=320\textrm{B}$ tokens used when training the 8B model independently, even though the total computational cost is the same (Figure 3, (Top) and (Middle)). Thus, it is possible that progressive training benefits unfairly from exposure to a larger volume of unique training data. To isolate this factor, we introduced a controlled setting denoted as Progressive+Fixed Data, in which total amount of unique data is capped at $320\textrm{B}$ tokens – the same amount used when independently training the 8B model. Once the 320B tokens were consumed during the training of the 8B model, the remaining 460320=140460-320=140B tokens are reused, i.e., the second epoch begins, as illustrated in Figure 3 (Bottom).
The result is shown in Table 2, in which 8B Progressive+Fixed Data achieves comparable performance. This confirms that the observed improvements from progressive training are not attributable to access to a greater quantity of unique data.